# Highly Undersampled MRI Reconstruction via a Single Posterior Sampling of Diffusion Models

### v1

@@ 3.4 Training Data Preparation and Network Training
(2) mGRE Brain Data A total of 44,063 complex-valued slices (image size: 256256) were retrospectively obtained from 47 subjects (Institutional ethics board approvals have been obtained for all MRI-QSM brain data used in this work), which were acquired at 3T (Discovery 750, GE Healthcare, Milwaukee, WI), with the following parameters: 8 unipolar echoes, first TE / TE = 3 ms / 3.3 ms, TR = 29.8 ms, FOV = 256256128, resolution = 1 mm isotropic. The image normalization and subsampled k-space simulation were made the same as the pipeline described for fastMRI brain dataset.
@@ 4.1 Evaluation Datasets
128 in distribution single-coil brain slices from fastMRI at acceleration factors (AF) of 8 and 12 (1D Gaussian subsampling, ACS = 12 and 6 lines, respectively).
Out-of-distribution (OOD) fastMRI knee dataset (156 single-coil images) at AF of 12 (Gaussian 2D).
@@
Retrospective multi-coil brain (78 images at AF=12) and multi-coil knee (103 images at AF=15) datasets with ground truths.
Two 12-coil prospectively undersampled brain and knee slices (AF = and ) randomly selected from fastMRI testing dataset.
A brain data (size: ) retrospectively undersampled at AF = 8 using sampling masks adopted in [15] to evaluate SSDM-MRI’s performance on advanced QSM acceleration task.
@@ 4.2 Performance Evaluation and Numerical Metrics
P-LORAKS | CCPC | ADMM-Net | U-net | Score-MRI | SSDM-MRI |
2.67 s | 221.45 s | 1.91 s | 0.36 s | 315.18 s | 0.45 s |
P-LORAKS and CSPC were run on a Intel i5-12400 CPU Processor, while the other deep learning method were evaluated on a NVIDIA 3060 GPU.
@@
SSDM-MRI was compared with the two traditional algorithms and three deep learning methods: (1) scan-specific P-LORAKS [49], (2) Compressed Sensing with Phase Cycling (CSPC) [40], (3) the unrolling-based ADMM-Net [17], (4) end-to-end U-net based method [16], and (5) DM-based Score-based MRI [26, 29] ( overall sampling steps) on the above described datasets of different acceleration factors and subsampling patterns. The deep learning methods were re-trained using the same datasets that were originally used for training the SSDM-MRI. The reconstruction times of different methods on a image of size are summarized in Table 1, indicating that the reconstruction speed of SSDM-MRI is much faster than that of Score-MRI, since only one reverse step is needed.
@@
Performance evaluation experiments are carried out as follows: (1) The proposed SSDM-MRI was first compared with other reconstruction methods using the single-coil fastMRI brain dataset. (2) Then, it was tested on OOD fastMRI knee images and the simulated pathological lesion brain image to test its generalization capability. (3) Next, SSDM-MRI’s performances were further validated on both the retrospective and prospective multi-coil fastMRI data. (4) In addition, SSDM-MRI’s performance was evaluated on one 8-echo GRE brain for QSM acceleration following [15], which could compare various methods at the whole brain level instead of a single 2D slice, and evaluate their capability of preserving latent susceptibility information. (5) One more simulated study was carried out to investigate the performance of SSDM-MRI against various acceleration factors (AF = 4, 6, 8, 10, 12, respectively). (6) Finally, the stability of SSDM-MRI was compared with Score-MRI by inspecting the generation of uncertainty maps.
P-LORAKS | ADMM-NET | U-net | Score-MRI | SSDM-MRI | |||
Single-coil knee (156 images) | Gaussian 2D | 25.94 ± 2.031 | 25.56 ± 2.074 | 27.94 ± 2.343 | 27.46 ± 2.265 | 29.95 ± 2.144 | 29.49± 2.213 |
×8 | 0.76 ± 0.047 | 0.76 ± 0.044 | 0.79 ± 0.043 | 0.79± 0.056 | 0.81 ± 0.071 | 0.82 ± 0.067 | |
Gaussian 1D | 21.74 ± 1.754 | 21.27 ± 1.337 | 27.56 ± 2.318 | 25.96 ± 2.866 | 28.19 ± 2.879 | 28.08 ± 2.311 | |
×12 | 0.56 ± 0.081 | 0.54 ± 0.076 | 0.77 ± 0.023 | 0.75 ± 0.085 | 0.76 ± 0.023 | 0.78± 0.043 | |
Single-coil brain (128 images) | Gaussian 1D | 25.34 ± 1.613 | 25.11 ± 1.431 | 32.93 ± 2.557 | 31.71 ± 2.808 | 32.77 ± 1.995 | 33.83 ± 2.311 |
×8 | 0.82 ± 0.049 | 0.82 ± 0.062 | 0.92 ± 0.017 | 0.91 ± 0.038 | 0.92 ± 0.024 | 0.93± 0.029 | |
Poisson 2D | 24.77± 1.983 | 24.54 ± 2.454 | 29.11 ± 2.353 | 28.98 ± 1.967 | 29.29 ± 2.348 | 30.19 ± 2.103 | |
×15 | 0.70 ± 0.061 | 0.69 ± 0.051 | 0.86 ± 0.037 | 0.84 ± 0.066 | 0.88 ± 0.038 | 0.89 ± 0.034 | |
Multi-coil brain (78 images) | Gaussian 1D | 21.45 ± 1.217 | 20.71 ± 1.146 | 28.15 ± 2.254 | 27.71 ± 2.365 | 29.59 ± 1.897 | 31.28 ± 2.314 |
×12 | 0.63 ± 0.084 | 0.63 ± 0.078 | 0.85 ± 0.024 | 0.83 ± 0.028 | 0.89 ± 0.047 | 0.90 ± 0.031 | |
Random 1D | 20.41 ± 1.943 | 19.85 ± 2.575 | 26.14 ± 2.154 | 25.42 ± 2.993 | 26.67 ± 2.079 | 27.66 ± 2.473 | |
×12 | 0.54 ± 0.074 | 0.53 ± 0.066 | 0.78 ± 0.061 | 0.77 ± 0.083 | 0.85 ± 0.053 | 0.86 ±0.041 | |
Multi-coil knee (103 images) | Poisson 2D | 23.67± 2.185 | 22.57 ± 2.553 | 28.31 ± 2.438 | 27.98 ± 2.567 | 29.35 ± 2.379 | 29.89 ± 2.132 |
×15 | 0.62 ± 0.072 | 0.63 ± 0.063 | 0.75 ± 0.087 | 0.74 ± 0.076 | 0.77 ± 0.073 | 0.78 ± 0.067 | |
Gaussian 1D | 20.31 ± 2.743 | 20.15 ± 2.677 | 26.71 ± 2.791 | 26.42 ± 2.990 | 27.97 ± 2.673 | 29.01 ± 2.510 | |
×12 | 0.55 ± 0.078 | 0.54 ± 0.076 | 0.73 ± 0.074 | 0.71 ± 0.083 | 0.78 ± 0.053 | 0.78 ± 0.057 |
@@
Peak Signal to Noise Ratio (PSNR) and Structural Similarity (SSIM) [50] were reported for quantitative assessments of different methods on magnitude images, while PSNR and XSIM (a QSM-specific SSIM variants) [51] were reported for the QSM acceleration task. In addition, voxel-wise linear regressions on QSM values of Substantia Nigra (SN) and Red Nucleus (RN) regions were carried out to compare different methods.
@@ 5.1 Evaluation on single-coil retrospective data
Fig. 2 compares SSDM-MRI with various MRI accelerating methods on two in-distribution fastMRI single-coil brain slices (AF = 12 and 8, respectively). Overall, SSDM-MRI showed the highest PSNR and SSIM on both datasets, and the minimum errors in the zoomed-in region. P-LORAKS and CSPC showed substantial residual artifacts (red arrows), which are absent in Score-MRI and SSDM-MRI. Meanwhile, SSDM-MRI preserved finer structures (green arrows) compared with other methods.
SSDM-MRI’s performance is evaluated on two OOD images (i.e., a fastMRI knee and a pathological brain) in Fig. 3. SSDM-MRI resulted in the minimum error maps and the best fine image details for the knee slice, as pointed out by the green arrow. For the pathological brain image, SSDM-MRI led to the best metrics among all methods, while ADMM-Net led to the minimum error map on the brain lesion.
@@ 5.2.1 Evaluation on retrospective data
The proposed SSDM-MRI is compared with P-LORAKS, CSPC, ADMM-Net, U-net, and Score-MRI on two retrospective fastMRI brain and knee images (12-coil) were compared in Fig. 4 at AF = 12 and 15, respectively. P-LORAKS and CSPC led to apparent artifacts and reconstruction errors at these high AFs, while U-net resulted in over-smoothed results. SSDM-MRI not only led to the best numerical metrics but also showed the best fine structures (green arrows), compared with other methods.
Table 2 summarizes the performances of different methods on four retrospective dataset obtained from fastMRI: (1) single-coil brain, (2) single-coil knee, (3) multi-coil brain, and (4) multi-coil knee images at different AFs and different types of sampling masks. SSDM-MRI achieved on average the best PSNR and SSIMs in nearly all situations, compared with the other algorithms.
@@ 5.2.2 Evaluation on prospective data
Different methods for two prospectively undersampled data are compared in Fig. 5. Similar to the simulation case, traditional algorithms, ie., P-LORAKS, CSPC demonstrated substantial artifacts (red arrows), which were successfully removed by ADMM-Net, U-net, Score-MRI, and SSDM-MRI. In addition, SSDM-MRI also significantly reduced the image blurring apparent in U-net and demonstrated the best fine structures as pointed out by the green arrows.
@@ 5.3 Comparison on QSM acceleration task
Fig. 6 compares different methods on a QSM brain volume at 8 acceleration. While all methods showed dramatic detail loss, compared to the fully-sampled QSM ground truth. SSDM-MRI resulted in the minimum error map and the best PSNR and XSIM (38.80 and 0.66). Linear regression was also performed on susceptibility values of various reconstruction methods against those of the ground truth, as shown Fig. 6(b). It confirmed that the proposed SSDM-MRI achieved the most accurate quantitative measurements in both SN ( =0.93) and RN ( =0.81) regions.
@@ 5.4 The effects of accelerating factors
The reconstructions of SSDM-MRI under 5 different AFs are shown in Fig. 7, on one retrospectively subsampled brain and a knee image. None of the reconstructed images showed apparent artifacts; however, as the AFs increases, more fine details gradually became blurred, which was also confirmed in the error maps.
@@ 5.5 Reconstruction stability of SSDM-MRI
Fig. 8 illustrates the uncertainty maps of SSDM-MRI and Score-MRI of 5 repetitive reconstructions on one knee slice at AF = 4 , 8, 12 respectively. As the AF increases from 4 to 12, both DM-based methods exhibited a notable increase in uncertainty images, with the Score-MRI demonstrating slightly larger standard deviations than the proposed SSDM-MRI.

### v2

@@ 3.5 Training Data Preparation and Network Training
(2) mGRE Brain Data A total of 44,063 complex-valued slices (image size: 256256) were retrospectively obtained from 47 subjects (Institutional ethics board approvals have been obtained for all MRI-QSM brain data used in this work), which were acquired at 3T (Discovery 750, GE Healthcare, Milwaukee, WI), with the following parameters: 8 unipolar echoes, first TE / TE = 3 ms / 3.3 ms, TR = 29.8 ms, FOV = 256 mm 256 mm 128 mm, resolution = 1 mm isotropic. The image normalization and subsampled k-space simulation were made the same as the pipeline described for fastMRI brain dataset.
@@ 4.1 Evaluation Datasets
300 in distribution single-coil brain slices from fastMRI at acceleration factors (AF) of 8 and 12 (1D Gaussian subsampling, ACS = 12 and 6 lines, respectively), and 15 (2D Poisson).
300 Out-of-distribution (OOD) fastMRI knee dataset at AF of 8 (1D Gaussian) and 12 (2D Gaussian).
@@
Retrospective 12-coil brain (320 images at AF = 12, 1D Gaussian and random mask) and 15-coil knee (320 images at AF = 12 of 1D Gaussian masks and 15 of 2D Poisson masks) datasets with ground truths.
Two 16-coil prospectively undersampled brain images (AF = and ) randomly selected from fastMRI test dataset.
A brain subject of five repetitive mGRE scans (size: ) acquired at different orientations, i.e., neutral head position (left 23°, right 17°, flexion 18°, and extension 21° tilt angles from the neutral head position) retrospectively undersampled at AF = 8 using sampling masks adopted in [20] to evaluate SSDM-MRI’s performance on advanced QSM acceleration task. COSMOS-QSM images were reconstructed based on the five repetitive scans of the mGRE brain subject at all five different head orientations, and the magnitude-based R2* images were calculated using the data of neutral head orientation.
@@ 4.2 Performance Evaluation and Numerical Metrics
SSDM-MRI was compared with the two traditional algorithms and four deep learning methods: (1) scan-specific P-LORAKS [11] (LORAKS type = “S”, rank = 40, and maximum iteration number = 100), (2) Compressed Sensing with Phase Cycling (CS) [10] ( = 0.003, = 0.005, inner iteration number = 10, and out iteration number = 100), (3) the unrolling-based MoDL[26] method, (4) end-to-end U-net based method [23], and (5) DM-based Score-based MRI [33] ( overall sampling steps), and (6) CCDF [40] (20 reverse steps as set in the original work) on the above described datasets of different acceleration factors and subsampling patterns. The data consistency operation was applied to all the methods involved in this work for fair comparisons. The deep learning methods were re-trained using the same datasets that were originally used for training the SSDM-MRI. The reconstruction times of different methods on a image of size are summarized in Table 1, indicating that the reconstruction speed of SSDM-MRI is much faster than that of Score-MRI, since only one reverse step is needed.
| P-LORAKS | CS | MoDL | U-net | Score-MRI | CCDF | SSDM-MRI |
| 2.67 s | 221.45 s | 1.91 s | 0.36 s | 315.18 s | 7.45 s | 0.45 s |
P-LORAKS and CS were run on a Intel i5-12400 CPU Processor, while the other deep learning method were evaluated on a NVIDIA 3060 GPU.
@@
Performance evaluation experiments are carried out as follows: (1) First, the proposed SSDM-MRI was compared with other reconstruction methods using the single-coil fastMRI brain dataset to demonstrate its superior performance on the in-distribution data. (2) Subsequently, it was tested on OOD fastMRI knee images and the simulated pathological lesion brain image to test its generalization capability. (3) Furthermore, SSDM-MRI’s performances were further validated on both the retrospective and prospective multi-coil fastMRI data to further demonstrate its generalization capability. (4) In addition, SSDM-MRI’s performance was evaluated on one 8-echo GRE brain subject for QSM acceleration following [20], which could compare various methods at the whole brain level instead of a single 2D slice, and evaluate their capability of preserving latent quantitative susceptibility information. This experiment can also compare the performance of SSDM-MRI with other methods from the perspective of MRI phase image reconstruction, which is often neglected in previous works. R2* image-based comparisons were also conducted to explore SSDM-MRI’s applicability in multi-echo magnitude imaging contexts. (5) One more simulated study was carried out to investigate the performance of SSDM-MRI against various acceleration factors (AF = 4, 6, 8, 10, 12, respectively). (6) Moreover, the stability of SSDM-MRI was compared with Score-MRI by inspecting the generation of uncertainty maps and a quantitative uncertainty comparison of the proposed SSDM-MRI with the other two DM methods, i.e., Score-MRI and CCDF. (7) Finally, an ablation study was performed to investigate the effectiveness of the proposed iterative selective distillation algorithm (paired with the proposed shortcut reverse sampling), the number of distillation iterations, and the data consistency (DC) module on the reconstruction result.
| Dataset | Mask | P-LORAKS | CS | MoDL | U-net | Score-MRI | CCDF | SSDM-MRI | ||||
Gaussian 1D | 25.29 1.612∗∗ | 25.16 1.439∗∗ | 32.87 2.563∗∗ | 31.74 2.802∗∗ | 32.82 1.987 | 0.018* | 32.94 2.144 | 0.015* | 33.97 1.974 | ||
| 8 | 0.81 0.048∗∗ | 0.82 0.061∗∗ | 0.91 0.018∗∗ | 0.90 0.039∗∗ | 0.92 0.025 | 0.020* | 0.91 0.031 | 0.016* | 0.94 0.021 | |||
| Poisson 2D | 24.80 1.981∗∗ | 24.49 2.461∗∗ | 29.15 2.360∗∗ | 28.94 1.973∗∗ | 29.23 2.340 | 0.016* | 29.11 2.414 | 0.013* | 30.42 2.144 | |||
| 15 | 0.71 0.060∗∗ | 0.68 0.052∗∗ | 0.86 0.038∗∗ | 0.85 0.065 ∗∗ | 0.87 0.037 | 0.018* | 0.87 0.039 | 0.015* | 0.90 0.031 | |||
Gaussian 1D | 25.90 2.038∗∗ | 25.61 2.069∗∗ | 27.88 2.336∗∗ | 27.50 2.259∗∗ | 29.81 2.152 | 0.089 | 29.63 2.378 | 0.062 | 29.90 2.121 | ||
| 8 | 0.77 0.046∗∗ | 0.75 0.043∗∗ | 0.78 0.042∗∗ | 0.79 0.057∗∗ | 0.81 0.072 | 0.017* | 0.81 0.083 | 0.014* | 0.83 0.067 | |||
| Gaussian 2D | 21.78 1.749∗∗ | 21.23 1.341∗∗ | 27.61 2.311∗∗ | 25.92 2.872∗∗ | 28.14 2.885 | 0.017* | 27.94 2.693 | 0.012* | 28.48 2.343 | |||
| 12 | 0.55 0.082∗∗ | 0.53 0.077∗∗ | 0.76 0.024∗∗ | 0.74 0.086∗∗ | 0.77 0.022 | 0.016* | 0.77 0.061 | 0.015* | 0.80 0.031 | |||
Gaussian 1D | 21.49 1.214∗∗ | 20.75 1.152∗∗ | 28.11 2.261∗∗ | 27.66 2.370∗∗ | 29.55 1.905 | 0.016* | 29.67 2.314 | 0.013* | 31.41 1.924 | ||
| 12 | 0.62 0.083∗∗ | 0.64 0.079∗∗ | 0.85 0.025∗∗ | 0.82 0.029∗∗ | 0.89 0.046 | 0.017* | 0.88 0.049 | 0.012* | 0.91 0.034 | |||
| Random 1D | 20.37 1.947∗∗ | 19.88 2.572∗∗ | 26.19 2.161∗∗ | 25.47 2.987∗∗ | 26.72 2.087 | 0.016* | 26.78 2.531 | 0.012* | 28.04 1.941 | |||
| 12 | 0.53 0.075∗∗ | 0.52 0.065∗∗ | 0.77 0.062∗∗ | 0.76 0.082∗∗ | 0.85 0.054 | 0.018* | 0.85 0.061 | 0.015* | 0.89 0.047 | |||
Poisson 2D | 23.72 2.181∗∗ | 22.52 2.547∗∗ | 28.27 2.444∗∗ | 28.01 2.560∗∗ | 29.30 2.385 | 0.016* | 29.13 2.247 | 0.014* | 30.17 2.148 | ||
| 15 | 0.63 0.071∗∗ | 0.62 0.062∗∗ | 0.74 0.088∗∗ | 0.73 0.077∗∗ | 0.77 0.072 | 0.017* | 0.76 0.079 | 0.013* | 0.80 0.069 | |||
| Gaussian 1D | 20.28 2.749∗∗ | 20.19 2.673∗∗ | 26.68 2.797∗∗ | 26.45 2.984∗∗ | 27.92 2.679 | 0.017* | 28.04 2.776 | 0.012* | 29.38 2.531 | |||
| 12 | 0.56 0.079∗∗ | 0.55 0.077∗∗ | 0.72 0.075∗∗ | 0.70 0.084∗∗ | 0.77 0.052 | 0.017* | 0.76 0.071 | 0.013* | 0.81 0.049 | |||
-values were calculated using t-tests for comparison between the proposed SSDM-MRI and the other methods in terms of PSNR and SSIM, respectively.
* indicates 0.05 and ** indicates 0.01 (note that -values 0.01 are not displayed to save more space for better visualization)
@@
Peak Signal to Noise Ratio (PSNR) and Structural Similarity (SSIM) [54] were reported for quantitative assessments of different methods on magnitude and R2* images, while PSNR, High Frequency Error Norm (HFEN) [55], and XSIM [56] (a specific SSIM implementation with optimization of internal parameters for evaluating QSM images, with hyperparameters set as follows: , , and ), were reported for the QSM acceleration task. In addition, voxel-wise linear regressions on QSM values of five deep grey matter regions, i.e., Substantia Nigra (SN), Red Nucleus (RN), Globus Pallidus (GP), Caudate (CN), and Putamen (PU), were carried out to compare different methods.
@@ 5.1 Evaluation on single-coil retrospective data
Fig. 3 compares SSDM-MRI with various MRI accelerating methods on two in-distribution fastMRI single-coil brain slices (AF = 12 and 8, respectively). Overall, SSDM-MRI showed the highest PSNR and SSIM on both datasets, and the minimum errors in the zoomed-in region. P-LORAKS and CS showed substantial residual artifacts (yellow arrows). All the other deep learning-based methods demonstrated over-smoothing and loss of fine details (red arrows) in the zoomed-in region, while the proposed SSDM-MRI better preserved the fine structures (green arrows).
SSDM-MRI’s performance is evaluated on two OOD images (i.e., a fastMRI knee image and a pathological brain slice) in Fig. 4. SSDM-MRI resulted in the minimum error maps and the best fine image details for the knee slice, as pointed out by the green arrow, while all other methods showed apparent reconstruction errors. For the pathological brain image, SSDM-MRI led to the best metrics (30.31 PSNR and 0.95 SSIM) among all methods with the minimum error map in the zoomed-in region, while P-LORAKS and CS showed substantial artifacts (yellow arrows).
@@ 5.2.1 Evaluation on retrospective data
The proposed SSDM-MRI is compared with P-LORAKS, CS, MoDL, U-net, and Score-MRI on two retrospective fastMRI brain (12 channels) and knee images (15 channels) in Fig. 5 at AF = 12 and 15, respectively. P-LORAKS and CS led to apparent artifacts and reconstruction errors at these high AFs, while MoDL, U-net, and CCDF resulted in over-smoothed results. SSDM-MRI not only led to the best numerical metrics but also showed the best fine structures (pointed out by the green arrows).
Table 2 summarizes the performances of different methods on four retrospective dataset obtained from fastMRI: (1) single-coil brain, (2) single-coil knee, (3) multi-coil brain, and (4) multi-coil knee images at different AFs and different types of sampling masks. SSDM-MRI achieved on average the best PSNR and SSIMs for all four datasets, compared to the other algorithms.
@@ 5.2.2 Evaluation on prospective data
Different methods for two prospectively undersampled data are compared in Fig. 6. Similar to the simulation case, traditional algorithms, i.e., P-LORAKS, CS demonstrated susbtantial aliasing artifacts (yellow arrows), which were absent in the other methods. SSDM-MRI significantly reduced the image blurring apparent in U-net, MoDL, Score-MRI and CCDF in the zoomed-in region and demonstrated the best image fine details (green arrows).
@@ 5.3 Comparison on QSM acceleration task
Fig. 7 compares different methods on an mGRE brain volume at 8 acceleration. The proposed SSDM-MRI resulted in the best phase-based COSMOS-QSM results and the magnitude-based R2* images with the minimum error maps and the best quantitative metrics. Specifically, SSDM-MRI achieved the best PSNR, XSIM, and HFEN (39.34, 0.79, and 0.276) among all QSM reconstructions. Linear regression was also performed on the susceptibility values of various reconstruction methods against those of the ground truth, as shown in Fig. 7(c), which confirmed that the proposed SSDM-MRI achieved the most accurate quantitative measurements in the five deep grey matter regions ( = 0.91, 0.76, 0.94, 0.93, and 0.88 for Substantia Nigra, Red Nucleus, Globus Pallidus, Caudate, and Putamen, respectively).
@@ 5.4 The effects of accelerating factors
The reconstructions of SSDM-MRI under 5 different AFs are shown in Fig. 8, on a retrospectively subsampled brain and a knee image. None of the reconstructed images showed apparent artifacts; however, as the AFs increases, more fine details gradually became blurred, which was also confirmed in the error maps.
@@ 5.5 Reconstruction stability of SSDM-MRI
Fig. 9 illustrates the uncertainty maps of SSDM-MRI and Score-MRI of 5 repetitive reconstructions on one knee slice at AF = 4, 8, 12, respectively. As the AF increases from 4 to 12, all DM-based methods exhibited a notable increase in uncertainty images, with the Score-MRI and CCDF demonstrating slightly larger standard deviations than the proposed SSDM-MRI, as highlighted by the red arrows. Furthermore, a quantitative uncertainty analysis was also reported in Table 3 by calculating the mean and standard deviations of SSDM-MRI, Score-MRI, and CCDF over 10 repetitive reconstructions on a fastMRI brain image at AF = 8. SSDM-MRI achieved higher mean PSNR/SSIM values with smaller standard deviations than Score-MRI and CCDF, suggesting more stable and robust reconstructions. For example, the PSNR standard deviation of SSDM-MRI is only 0.355, compared to 0.512 and 0.634 for Score-MRI and CCDF, respectively.
5.6 Ablation study on data consistency, the proposed selective distillation algorithm, and the number of distillations
| Method | PSNR | SSIM |
| Score-MRI | ||
| CCDF | ||
| SSDM-MRI (proposed) |
An ablation study was conducted to investigate the influences of data consistency (DC), the proposed selective distillation (SD) algorithm, and the number of distillations on the SSDM-MRI’s performance, as shown in Fig. 10. Overall, the performance of the proposed SSDM-MRI slightly decreases with increasing number of distillations, from the original 28.75 PSNR and 0.923 SSIM (128 reverse steps) to 28.32 PSNR and 0.906 SSIM (only 1 reverse step) after 4 distillation steps, which are still much better than the corresponding 1-step result before distillation (9.52/0.152 of PSNR/SSIM). Both DC operations and the proposed SD algorithms are important for SSDM-MRI to achieve high-quality reconstructions after distillation. For instance, distilling the pretrained conditional DM 4 times using the proposed iterative selective distillation with the incorporation of DC only observes a decrease of 0.43 and 0.017 in PSNR and SSIM, respectively, compared to a 1.78/0.047 PSNR/SSIM decrease for SSDM-MRI without Selective Distillation (i.e., SSDM-MRI using conventional progressive distillation), and a 1.39/0.044 PSNR/SSIM decrease for SSDM-MRI without data consistency. This improvement in alleviating the performance decrease after distillation might be due to the reason that the proposed iterative selective distillation algorithm focuses only on the most important parts of the reverse sampling, rather than the entire reverse path from the pure noise to clean images.